OpenStudy (saaarrrahhhh):

Help me understand exponential functions?? The graph of an exponential function of the form y = f(x) = a^x passes through the points (0, ?) and (1, ?) what are the y coordinates??

OpenStudy (julisam):

Thank you.

OpenStudy (latinc):

f(x) = y; (x,y) = (0,?) f(0) = a^0 ----> f(0) = 1 (x,y) = (1,?) f(1) = a^1 ------> f(1) = a

OpenStudy (latinc):

did that help?

OpenStudy (saaarrrahhhh):

yes!! thank you @latinC